Children’s hospital predicts variable costs of 70% of total revenue and fixed costs of $42 million per year for the coming year.

  1. Compute the break-even point expressed in total revenue.
  2. Children’s Hospital expects total revenue of $150 million from 200,000 patient days. Compute expected profit (a) if costs behave as expected, and (b) if variable costs are 10% greater than predicted.
Answer #1

Answer

--Break even point = Fixed Cost / CM ratio = $42 million / (100% - 70%) = $ 42 million / 30% = $ 140 millions Answer

--Expected Profits:
(a) = ($ 150 million sales x 30%) - $ 42 million fixed cost = 45 - 42 = $ 3 millions Net Income Answer

(b) = (150 millions sales x 20%) - $ 42 millions = 30 - 42 = $ 12 million Net Loss Answer